Paul and Malachi were in the process of modifying the power supply to keep the fan running continuously and trying to eliminate the disruptions in the voltage. We expected we could do it while the power supply was still running. I guess we were wrong, because it started sparking and arcing and smoking. And as everybody knows, once you release the magic smoke from inside electrical components, the components don't work, anymore.
In the photo below, I am pointing my web cam at my desk towards the damaged power supply so Mathieu in France can see the black area.
So we swapped out the power supply with another one with an improved, temperature compensated interface, made the fan run continuously from the external 9V power supply that powers HUGNetLab and the pressure sensor, and we were back on line in an hour. When the experiment comes back on-line we should see much cleaner data without all the voltage ripples.
Experiment: Next step
As long as our current experiment was disrupted, we are moving forward to another phase of the experiment in which we attempt to reload the wire with hydrogen. Steps towards this:
- Let it cool to ambient,
- Pull a vacuum in the cell
- Pressurize to 3.5 bar H2
- Heat to approximately 170 T_Mica. Directly in one step, not ramped, or smaller, stepped increases. But some fine tuning may be necessary.
Our thinking is that the wire has become unloaded at the higher temperature and current as indicated by the increase in impedance over the last couple days. We will attempt to reload it and then attempt triggering again.
Addition: It is complete and re-warming before I managed to complete this blog entry.
Better Control of Room Temp:
One thing that has been troublesome since the weather got colder. We tried putting electric space heaters with thermostats into the room, but that causes cycling of temperature. So the space heaters were not cutting it. Instead, we are working on a much tighter, proportional control system with feedback. Assembly of the system is in process right now and it should start working tomorrow.
Euro Cell Report:
Mathieu has been kinda quiet on the blog, but has been making good progress in the background. He reports completing 3 calibration runs in Helium. At this point, I believe that will be considered just practice, because he discovered something that he is working to change.
When we put together the Euro Cell, we had glass that was slightly rough on the ends. It was not sealing too well. So we added high temperature silicone to the rough glass ends to make a better seal. It worked. But Mathieu discovered it had a bad side effect. After hitting the very high temperatures, he found red silicone on the inside of the cell coating the wires, the mica, and everything.
So Mathieu is taking the opportunity to replace the mica with some 0.5 mm mica, instead.
"I used a massicot (not the guillotine kind ;) paper-cutter to cut the 0.5mm thick mica then smoothed with some sand paper. Works well!"- Mathieu
Of course, this will change the cell characteristics some, so he will have to begin the calibrations again, but he would have had to, anyway because the silicone contamination would have altered the results.
Visitors: LENR Tourism in Minnesota
We had a couple of LENR enthusiasts pop by to visit the test cell. Very nice gentlemen from the Minneapolis area. They were very enthusiastic and enjoyed looking at the cell they had been reading about and following so closely. I wonder if we will see more LENR tourists soon?
We are updating the viewer, again, to work with 30 second data. We will post about this more as it gets released.